OpenStudy (anonymous):

A culture of bacteria started with 3,000 after 2 hours it grew to 3300. How much will grow at 16 hours? Round to nearest whole. P=Ae^kt

OpenStudy (unklerhaukus):

if t=0 P=A ie A=3,000

OpenStudy (unklerhaukus):

P=3300 when t=2hours

OpenStudy (unklerhaukus):

you still need to find k,

OpenStudy (anonymous):

=(

OpenStudy (unklerhaukus):

\[P=Ae^{kt}\]\[3300=3000e^{2k}\]\[\frac{3300}{3000}=e^{2k}\]\[11/10=e^{2k}\]\[\ln{11/10}=2k\]\[k=\frac{\ln(\frac{11}{10})}{2} \]

OpenStudy (unklerhaukus):

so \[P(t)=Ae^{kt}\rightarrow P(t)=3000e^{\frac{\ln(11/10)}{2}t}\] to find P at t=16 \[P(16)=3000e^{\frac{\ln(11/10)}{2}16}\]